# A localisation phase transition for the catalytic branching random walk

### v1

@@ 1.4 Discussion of the literature
The fact that $N_{t}$ is square integrable for all $t\geq 0$ follows from the fact that $(N_{t})_{t\geq 0}$ is stochastically dominated by the total population of a binary branching process with reproduction rate $\lambda_{0}$, whose square is well-known to be integrable. For all $0\leq s\leq t$, we let $f_{s}(t)=\mathbb{E}[M_{t}|\mathcal{F}_{s}]$, where $(\mathcal{F}_{s})_{s\geq 0}$ is the natural filtration of the process $(\Pi_{s})_{s\geq 0}$. It is standard to see, using the definition of the process $(\Pi_{t})_{t\geq 0}$ as a Markov jump process, that, for all $s\geq 0$, $t\mapsto f_{s}(t)$ is continuous on $[s,+\infty)$, and moreover, that for all $t>s$, $\frac{\mathrm{d}}{\mathrm{d}t}f_{s}(t)=0$, from which we deduce that $(M_{t})_{t\geq 0}$ is indeed a martingale. To see that it is bounded in $L^{2}$, note that for any fixed $t\geq 0$,
[…]
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[M_{t}^{2}]=\mathbb{E}[M_{t}\rho_{t}\mathrm{e}^{-\int_{0}^{t}\rho_{s}\,\mathrm{d}s}]\leq\lambda_{0}\mathbb{E}[M_{t}^{2}]^{1/2}\mathrm{e}^{-\lambda t},$
which implies
$\sup_{t\geq 0}\mathbb{E}[M_{t}^{2}]\leq\Big{(}1+\frac{\lambda_{0}}{\lambda}\Big{)}^{2},$
@@
For the next step we need some additional notation: For all $x\in\mathbb{Z}^{d}$, we let $\Pi_{t}^{x}$ be the empirical measure of the particles in the catalytic branching random walk starting with one particle at site $x$. Given any measure $\mu$ on $\mathbb{Z}^{d}$, we let $\Pi_{t}\cdot\mu=\sum_{x\in\mathbb{Z}^{d}}\mu(x)\Pi_{t}^{x}$.
@@
If there exists a probability measure $\nu$ such that $\lim_{t\to\infty}\widehat{\Pi}_{t}\to\nu$, almost surely for the topology of weak convergence, then $\mathbb{E}[\Pi_{t}\cdot\nu]\cdot\mathrm{e}^{-\rho t}=\nu$, for all $t\geq 0$, with $\rho=\lambda+(\lambda_{0}-\lambda)\nu_{0}$.
@@
because $\mathbb{E}[M_{\infty}]=1$. Furthermore, because by assumption $\rho_{t}\to\rho$ almost surely as $t\uparrow\infty$, we get $\int_{s}^{t+s}\rho_{u}\,\mathrm{d}u=\rho t+o(1)$ as $s\uparrow\infty$. Hence,
@@
Also, by Cauchy-Schwarz’s inequality, for all $y\in\mathbb{Z}^{d}$,
@@
Moreover, when $\|y\|>2\|x\|$, we can bound the probability on the right-hand side by the probability that a simple random starting from $y$ enters the ball $B(0,\|x\|)$. Thus there exist two positive constants $c$ and $C$ such that, for all $y\in\mathbb{Z}^{d}$ satisfying $\|y\|>2\|x\|$,
$\mathbb{P}(\Pi_{t}^{y}(x)\neq 0)\leq C\exp(-c\|y\|^{2}/t).$
@@
Now, let $\nu$ be a probability measure on $\mathbb{Z}^{d}$ satisfying (1.1). To simplify notation, we set $\varepsilon=\lambda_{0}-\lambda$. Consider a simple random walk $(S_{n})_{n\geq 0}$ on $\mathbb{Z}^{d}$, and let $\tau_{0}$ be the first return time to 00, i.e. $\tau_{0}=\inf\{n\geq 1:S_{n}=0\}$. Equation (1.1) shows that for any $x\neq 0$, on the event that $S_{0}=x$, the process $(M_{n})_{n\geq 0}$, defined for $n\geq 0$ by
@@
11 11 $=\nu_{0}+\nu_{0}\sum_{k=1}^{\infty}\frac{1}{(1+\varepsilon\nu_{0})^{k}}\cdot\mathbb{P}(0\notin\{S_{1},\dots,S_{k}\})$
@@
As proved in Section 3.1, any probability measures $\nu$ satisfying (1.1) satisfies $\nu_{0}>0$. For any positive real number $\nu_{0}$, we use (3.1) to define a positive measure $\nu$ on $\mathbb{Z}^{d}$ that satisfies (1.1), by the Markov property. Now note that a simple random walk on $\mathbb{Z}^{d}$ started at some $x$ cannot reach 00 before time $\|x\|_{\infty}$. Thus, $\nu_{x}\leq\nu_{0}(1+\varepsilon\nu_{0})^{-\|x\|_{\infty}}$, which is summable. Hence it only remains to show that, for all $\varepsilon>\gamma_{d}$, there exists a unique value of $\nu_{0}$ making $\nu$ a probability measure. This is equivalent to showing that there exists a unique value of $\nu_{0}$ such that (3.1) is satisfied. Letting $u=\varepsilon\nu_{0}$, we see that this is equivalent to the existence and uniqueness of a solution to the equation $f(u)=\varepsilon$, where $\varepsilon>\gamma_{d}$ is fixed and for $u>0$,
@@
(Note that this last expression of $f$ could also have been derived directly from (1.2) and (3.1).) In particular $f$ is strictly increasing on $(0,\infty)$, converges to +$\infty$+ ∞, as $u$ goes to infinity, and by dominated convergence, we can see that it converges to $\gamma_{d}$ as $u$ goes to zero, which proves well that for any $\varepsilon>\gamma_{d}$, the equation $f(u)=\varepsilon$ has a unique solution.
@@
$f(u)=(1+u)\cdot\big{(}1-G(\frac{1}{1+u})\big{)},$
where $G(z)=\mathbb{E}[z^{\tau_{0}}]$, is the generating function of $\tau_{0}$. Now for $n\geq 1$, denote by $\tau_{n}$ the $n$-th return time to the origin of the walk, and let $G_{n}(z)=\mathbb{E}[z^{\tau_{n}}]$. By independence between the successive return times to the origin, one has $G_{n}(z)=G(z)^{n}$, for any $n\geq 1$. In particular one has for any $10<z<$ 1,
@@
[…]
where $p_{t}(0,0)$ is the probability that a continuous time simple random walk is at the origin at time $t$. Since by definition the time between any consecutive jumps is a mean one exponential random variable, one also has
[…]
@@
$G(\frac{1}{1+\rho})=\frac{\rho}{1+\rho}.$
In particular we recover well the same equation for $\rho$ as our equation defining $\nu_{0}$, namely $f(\nu_{0})=1$, which agrees with the fact that in our setting, when $\lambda_{0}=1$, and in the limit $\lambda\to 0$, the exponential growth rate is given by $\nu_{0}$.
@@ 4.1 Measure-valued Pólya processes
where, given $m_{n}$, $\xi(n+1)$ is a random variable of distribution $m_{n}P/m_{n}P(E)$, and where, given $\xi(n+1)$, $R^{\scriptscriptstyle(n+1)}_{\xi(n+1)}$ is an independent copy of $R^{\scriptscriptstyle(1)}_{\xi(n+1)}$.
@@
To prove localisation, we use [MV20, Theorem 1] (see Section 1.4 where the case of $R^{\scriptscriptstyle(1)}$ being a signed-measure is discussed). We prove that $(m_{n})_{n\geq 0}$ satisfies the assumptions (T), (A1), (A’2), (A3) and (A4) of [MV20], which, for convenience, we copy here: first define $R=\mathbb{E}[R^{\scriptscriptstyle(1)}]$, meaning that, for all $x\in E$, for all measurable sets $A\subset E$, $R_{x}(A)=\mathbb{E}[R^{\scriptscriptstyle(1)}_{x}(A)]$. We also define, for all $x\in E$,
@@
  • (A1)
    There exists $c>0$ and $\beta>1$ such that, for all $x\in E$, almost surely,
    $c\leq\inf_{x\in E}Q_{x}(E)\leq\sup_{x\in E}Q_{x}(E)\leq 1\,\text{ and }\,\sup_{x\in E}\,\mathbb{E}\big{|}Q_{x}^{\scriptscriptstyle(1)}(E)-Q_{x}(E)\big{|}^{\beta}<+\infty.$ (4.3)
@@
    the continuous-time pure jump Markov process $X$ with sub-Markovian jump kernel $Q-I$ admits a quasi-stationary distribution $\nu\in\mathcal{P}(E)$ (the set of all probability measures on $E$). We further assume that the convergence of $\mathbb{P}_{\alpha}(X_{t}\in\cdot\,|\,X_{t}\neq\partial)$ holds uniformly with respect to the total variation norm on a neighbourhood of $\nu$ in $\mathcal{P}_{C}(E)=\{\alpha\in\mathcal{P}(E)\mid\alpha\cdot V^{1/q}\leq C\}$, for each $C>0$, where $q=p/(p-1)$.
@@
    for all bounded continuous functions $f:E\rightarrow\mathbb{R}$, $x\in E\mapsto R_{x}f$ and $x\in E\mapsto Q_{x}f$ are continuous.
@@
(A1) above is not as in [MV20], but this version of (A1) suffices, as discussed in [MV20, Section 1.5] (see Equation (3) therein).
@@
Under Assumptions (T), (A1), (A’2), (A’3), and (A4), if $m_{0}\cdot V<\infty$ and $m_{0}P\cdot V<\infty$, then the sequence of random measures $(m_{n}/n)_{n\geq 0}$ converges almost surely to $\nu R$ with respect to the topology of weak convergence. Furthermore, if $\nu R(\mathbb{Z}^{d})>0$, then $(\tilde{m}_{n}:=m_{n}/m_{n}(\mathbb{Z}^{d}))_{n\in\mathbb{N}}$ converges almost surely to $\nu R/\nu R(\mathbb{Z}^{d})$ with respect to the topology of weak convergence.
@@
We also let $\tilde{\eta}_{0}=0$ and, for all $n\geq 1$, $\tilde{\eta}_{n}=\frac{\eta_{n}}{n}$. Finally, we set $\tau_{n}=\gamma_{1}+\cdots+\gamma_{n}$, where $\gamma_{n}=\frac{1}{n\tilde{\eta}_{n-1}Q(E)}$ for all $n\geq 1$. The proof of [MV20] relies on the fact that $(\tilde{\eta}_{n})_{n\geq 0}$ is a stochastic approximation: indeed, for all $n\geq 0$, one can write
@@
where $F(\mu)=\mu Q-\mu Q(E)\mu$ and $U_{n+1}=\tilde{\eta}_{n}Q(E)\delta_{\xi(n+1)}-\tilde{\eta}_{n}Q$. Remark 12 in [MV20] states that, without Assumption (A3), one can still conclude that $(\tilde{\eta}_{t})_{t\geq 0}$ is an asymptotic pseudo-trajectory in $\mathcal{P}_{C}(E)$ for the semi-flow induced in $\mathcal{P}_{C}(E)$ by the well-posed dynamical system
@@
(for all bounded functions $f=E\mapsto\mathbb{R}$). Now, by Theorem 3.7 in [Ben06], the limit set of a pre-compact asymptotic pseudo-trajectory is internally chain transitive and thus (by [Ben06, Proposition 5.3]) it does not contain any proper attractor. Now let $(X_{t})_{t\geq 0}$ be the continuous-time pure jump Markov process with sub-Markovian jump kernel $Q-I$; if, for all $t\geq 0$, we let $\mu_{t}$ be the distribution of $X_{t}$ conditioned on having survived until time $t$, then $(\mu_{t})_{t\geq 0}$ is a solution of (4.4). By Assumption (A3), we get that $\nu$ is an attractor for the flow defined by (4.4) on $\mathcal{P}_{C}(E)$. Also, all trajectories started in $\mathcal{P}_{C}(E)$ converge to $\nu$. Thus, the limit set of $(\tilde{\eta}_{t})_{t\geq 0}$ is $\{\nu\}$. Therefore, $\tilde{\eta}_{n}\to\nu$ almost surely as $n\uparrow\infty$, for the topology of weak convergence. One can then follow the rest of the arguments in [MV20] to conclude the proof. ∎
4.2 Checking Assumption (A1)

### v2

@@ 2.2 Proof of Part (ii) (weak delocalisation result)
The fact that $N_{t}$ is square integrable for all $t\geq 0$ follows from the fact that $(N_{t})_{t\geq 0}$ is stochastically dominated by the total population of a binary branching process with reproduction rate $\lambda_{0}$, whose square is well-known to be integrable. For all $0\leq s\leq t$, we let $f_{s}(t)=\mathbb{E}[M_{t}|\mathcal{F}_{s}]$, where $(\mathcal{F}_{s})_{s\geq 0}$ is the natural filtration of the process $(\Pi_{s})_{s\geq 0}$. It is standard to see, using the definition of the process $(\Pi_{t})_{t\geq 0}$ as a Markov jump process, that, for all $s\geq 0$, $t\mapsto f_{s}(t)$ is continuous on $s,+\infty)$, and moreover, that for all $t>s$, $\frac{\mathrm{d}}{\mathrm{d}t}f_{s}(t)=0$, from which we deduce that $(M_{t})_{t\geq 0}$ is indeed a martingale. To see that it is bounded in $L^{2}$, note that for any fixed $t\geq 0$,
$\mathbb{E}[N_{t+h}^{2}\mathrm{e}^{-2\int_{0}^{t}\rho_{s}\,ds}]$ $=h\mathbb{E}[\rho_{t}M_{t}(N_{t}+1)^{2}\mathrm{e}^{-\int_{0}^{t}\rho_{s}\,\mathrm{d}s}]+\mathbb{E}[(1-h\rho_{t}N_{t})M_{t}^{2}]+\mathcal{O}(h^{2})$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[M_{t}^{2}]=\mathbb{E}[M_{t}\rho_{t}\mathrm{e}^{-\int_{0}^{t}\rho_{s}\,\mathrm{d}s}]\leq\lambda_{0}\mathbb{E}[M_{t}^{2}]^{1/2}\mathrm{e}^{-\lambda t}.$
In other words,
$\frac{\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[M_{t}^{2}]}{2\sqrt{\mathbb{E}[M_{t}^{2}]}}\leq\frac{\lambda_{0}}{2}\mathrm{e}^{-\lambda t}.$
By integration, and because $\mathbb{E}[M_{0}^{2}]=1$, we get
$\sqrt{\mathbb{E}[M_{t}^{2}]}\leq 1+\frac{\lambda_{0}}{2\lambda}.$
Taking the square and then the supremum on $t$, we get
$\sup_{t\geq 0}\mathbb{E}[M_{t}^{2}]\leq\Big{(}1+\frac{\lambda_{0}}{2\lambda}\Big{)}^{2},$
@@
For the next step we need some additional notation: For all $x\in\mathbb{Z}^{d}$, we let $\Pi_{t}^{x}$ and $N_{t}^{x}$ be respectively the empirical measure of the particles and the total number of particles at time $t$ in the catalytic branching random walk starting with one particle at site $x$. Given any measure $\mu$ on $\mathbb{Z}^{d}$, we let $\Pi_{t}\cdot\mu=\sum_{x\in\mathbb{Z}^{d}}\mu(x)\Pi_{t}^{x}$.
@@
If there exists a probability measure $\nu$ such that $\lim_{t\to\infty}\widehat{\Pi}_{t}=\nu$, almost surely for the topology of weak convergence, then $\mathbb{E}[\Pi_{t}\cdot\nu]\cdot\mathrm{e}^{-\rho t}=\nu$, for all $t\geq 0$, with $\rho=\lambda+(\lambda_{0}-\lambda)\nu_{0}$.
@@
because $\mathbb{E}[M_{\infty}]=1$. Furthermore, because by assumption $\rho_{u}\to\rho$ almost surely as $u\uparrow\infty$, we get $\int_{s}^{t+s}\rho_{u}\,\mathrm{d}u=\rho t+o(1)$ as $s\uparrow\infty$. Since for any fixed $t>0$, $\int_{s}^{t+s}\rho_{u}\,\mathrm{d}u$ is uniformly bounded in $s$, and using that the martingale $(M_{t})_{t\geq 0}$ is bounded in $L^{2}$ by Lemma 3.1, we deduce by dominated convergence, that
@@
Also, by Cauchy-Schwarz’s inequality, for all $y\in\mathbb{Z}^{d}$ and $t\geq 0$,
@@
Moreover, when $\|y\|>2\|x\|$, we can bound the probability on the right-hand side using that the probability for a simple random starting from $y$ to enter the ball $B(0,\|x\|)$ by time $t$ is bounded by $C\exp(-c\|y\|^{2}/t)$, for some positive constants $c$ and $C$. -Hence a union bound on the set of particles alive at time $t$ ensures that, for all $y\in\mathbb{Z}^{d}$ satisfying $\|y\|>2\|x\|$,
$\mathbb{P}(\Pi_{t}^{y}(x)\neq 0)\leq C\mathrm{e}^{\lambda t}\cdot\exp(-c\|y\|^{2}/t).$
@@
Now, let $\nu$ be a probability measure on $\mathbb{Z}^{d}$ satisfying (1.1). To simplify notation, we set $\varepsilon=\lambda_{0}-\lambda$. Consider a simple random walk $(S_{n})_{n\geq 0}$ on $\mathbb{Z}^{d}$, and let $\tau_{0}$ be the first return time to 0, i.e. $\tau_{0}=\inf\{n\geq 1:S_{n}=0\}$. Equation (1.1) shows that for any $x\neq 0$, on the event that $S_{0}=x$, the process $(M_{n})_{n\geq 0}$, defined for $n\geq 0$ by
@@
$1$ $=\nu_{0}+\nu_{0}\sum_{k=1}^{\infty}\frac{1}{(1+\varepsilon\nu_{0})^{k}}\cdot\mathbb{P}(0\notin\{S_{1},\dots,S_{k}\})$
@@
As proved in Section 4.1, any probability measures $\nu$ satisfying (1.1) satisfies $\nu_{0}>0$. For any positive real number $\nu_{0}$, we use (4.1) to define a positive measure $\nu$ on $\mathbb{Z}^{d}$ that satisfies (1.1), by the Markov property. Now note that a simple random walk on $\mathbb{Z}^{d}$ started at some $x$ cannot reach 0 before time $\|x\|_{\infty}$. Thus, $\nu_{x}\leq\nu_{0}(1+\varepsilon\nu_{0})^{-\|x\|_{\infty}}$, which is summable. Hence it only remains to show that, for all $\varepsilon>\gamma_{d}$, there exists a unique value of $\nu_{0}$ making $\nu$ a probability measure. This is equivalent to showing that there exists a unique value of $\nu_{0}$ such that (4.1) is satisfied. Letting $u=\varepsilon\nu_{0}$, we see that this is equivalent to the existence and uniqueness of a solution to the equation $f(u)=\varepsilon$, where $\varepsilon>\gamma_{d}$ is fixed and for $u>0$,
@@
(Note that this last expression of $f$ could also have been derived directly from (1.2) and (4.1).) In particular $f$ is strictly increasing on $(0,\infty)$, converges to $\infty$, as $u$ goes to infinity, and by dominated convergence, we can see that it converges to $\gamma_{d}$ as $u$ goes to zero, which proves well that for any $\varepsilon>\gamma_{d}$, the equation $f(u)=\varepsilon$ has a unique solution.
@@
$f(u)=(1+u)\cdot\Big{(}1-G\Big{(}\frac{1}{1+u}\Big{)}\Big{)},$
where $G(z)=\mathbb{E}[z^{\tau_{0}}]$, is the generating function of $\tau_{0}$. Now for $n\geq 0$, denote by $\tau_{n}$ the $(n+1)$-th return time to the origin of the walk, and let $G_{n}(z)=\mathbb{E}[z^{\tau_{n}}]$. By independence between the successive return times to the origin, one has $G_{n}(z)=G(z)^{n+1}$, for any $n\geq 0$. In particular one has for any $0<z<1$,
@@
$\int_{0}^{\infty}\mathrm{e}^{-\rho t}\,p_{t}(0,0)\,dt=1,$ (4.5)
where $p_{t}(0,0)$ is the probability that a continuous time simple random walk is at the origin at time $t$. Since by definition the time between any two consecutive jumps is a mean one exponential random variable, one also has, with $\widetilde{\tau}_{n}$ the $n$-th return time to the origin for the continuous time random walk,
$\int_{0}^{\infty}\mathrm{e}^{-\rho t}\,p_{t}(0,0)\,dt$ $=\frac{1}{1+\rho}+\sum_{n\geq 1}\int_{0}^{\infty}\mathrm{e}^{-\rho t}\cdot\mathbb{E}\big{[}\mathrm{e}^{-(t-\widetilde{\tau}_{n})}\cdot\mathbf{1}\{\widetilde{\tau}_{n}\leq t\}\big{]}\,dt$
$=\frac{1}{1+\rho}+\frac{1}{1+\rho}\sum_{n\geq 1}\mathbb{E}[\mathrm{e}^{-\rho\widetilde{\tau}_{n}}]$
$=\frac{1}{1+\rho}+\frac{1}{1+\rho}\sum_{n\geq 0}G_{n}\Big{(}\frac{1}{1+\rho}\Big{)}.$
@@
$G\Big{(}\frac{1}{1+\rho}\Big{)}=\frac{\rho}{1+\rho}.$
In particular we recover well the same equation for $\rho$ as our equation defining $\nu_{0}$, namely $f(\nu_{0})=1$, which agrees with the fact that in our setting, when $\lambda_{0}=1$, and in the limit $\lambda\to 0$, the exponential growth rate is given by $\nu_{0}$, as shown by combining Lemma 3.1 and Theorem 1.1.
@@ 5.1 Measure-valued Pólya processes
where, given $m_{n}$, $\xi(n+1)$ is a random variable of distribution $m_{n}P/m_{n}P(E)$ with
$m_{n}P=\int_{x\in E}P_{x}\mathrm{d}m_{n}(x),$
and where, given $\xi(n+1)$, $R^{\scriptscriptstyle(n+1)}_{\xi(n+1)}$ is an independent copy of $R^{\scriptscriptstyle(1)}_{\xi(n+1)}$.
@@
To prove localisation, we use [MV20, Theorem 1] (see Section 1.4 where the case of $R^{\scriptscriptstyle(1)}$ being a signed-measure is discussed). We prove that $(m_{n})_{n\geq 0}$ satisfies the assumptions (T), (A’1), (A’2), (A’3) and (A4) below, which are a slight modification of the assumptions in [MV20] (we replace (A1) and (A3) therein by two other assumptions called (A’1) and (A’3)). First define $R=\mathbb{E}[R^{\scriptscriptstyle(1)}]$, meaning that, for all $x\in E$, for all measurable sets $A\subset E$,
$R_{x}(A)=\mathbb{E}[R^{\scriptscriptstyle(1)}_{x}(A)].$
We also define, for all $x\in E$,
@@
  • (A’1)
    For all $x\in E$, $Q_{x}(E)\leq 1$. Furthermore, there exists $c>0$ such that
    $\liminf_{n\uparrow\infty}\frac{m_{n}P(E)}{n}=\liminf_{n\uparrow\infty}\frac{1}{n}\sum_{i=1}^{n}Q^{\scriptscriptstyle(i)}_{\xi(i)}(E)\geq c\quad\text{ and }\quad\liminf_{n\uparrow\infty}\frac{1}{n}\sum_{i=1}^{n}Q_{\xi(i)}(E)\geq c.$
@@
    the continuous-time pure jump Markov process $X$ with sub-Markovian jump kernel QIQ-I admits a quasi-stationary distribution $\nu\in\mathcal{P}(E)$ (the set of all probability measures on $E$). Also, the convergence of $\mathbb{P}_{\alpha}(X_{t}\in\cdot\,|\,X_{t}\neq\partial)$ holds uniformly with respect to the total variation norm on a neighbourhood of $\nu$ in $\mathcal{P}_{C}(E)=\{\alpha\in\mathcal{P}(E)\mid\alpha\cdot V^{1/q}\leq C\}$, for each $C>0$, where q=p/(p1)q=p/(p-1). Finally, for all $\alpha\in\mathcal{P}_{C}(E)$, $\mathbb{P}_{\alpha}(X_{t}\neq\partial)\geq\mathrm{e}^{-(1-c_{0})t}$ for some constant $c_{0}>\theta$.
@@
    for all bounded continuous functions $f:E\rightarrow\mathbb{R}$, $x\in E\mapsto R_{x}\cdot f$ and $x\in E\mapsto Q_{x}\cdot f$ are continuous.
@@
We have replaced Assumption (A1) in [MV20] by Assumption (A’1). The fact that this can be done can be checked by going through the proofs in [MV20], which adapt straightforwardly. Indeed, here is a list of the changes one needs to make:
  • In [MV20, Lemma 3], one needs to change the definition of $\sigma_{k}$ to
    $\sigma_{k}=\inf\{n\geq k\colon m_{n}P(E)<c^{\prime}n\text{ or }\sum_{i=1}^{n}Q_{\xi(i)}(E)<c^{\prime}n\},$
    where $c^{\prime}\in(\theta,c)$. The proof of Lemma 3 then directly follows from (A’1).
  • Assumption (A1) is then used three times in the proof of [MV20, Lemma 5]; the same three arguments go through by using the fact that on $\{\ell\leq\sigma_{k}\}$, $\tilde{\eta}_{\ell-1}Q(E):=\sum_{i=1}^{\ell-1}Q_{\xi(i)}(E)\geq c^{\prime}(\ell-1)$ (instead of $\tilde{\eta}_{\ell-1}Q(E)\geq c_{1}=c$, where $c_{1}$ is the notation used in [MV20] for our constant $c$).
  • Assumption (A1) is also used in the proof of [MV20, Lemma 6]. There, one has to first fix an integer $k\geq 0$ and prove that there exists a random value of $C=C(k)$ such that $\tilde{\eta}_{n}\in\mathcal{P}_{C}(E)$ (the set of all probability measures $\mu$ on $E$ such that $\mu\cdot V^{\nicefrac{{1}}{{q}}}\leq C$) for all $n\leq\sigma_{k}$. Instead of using Assumption (A1), we use the fact that $\tilde{\eta}_{n}(E)\leq c^{\prime}n$ for all $n\leq\sigma_{k}$; with this modification, the proof goes through replacing $c_{1}$ (=c=c in this paper) by $c^{\prime}$ and only considering $n$ up to $\sigma_{k}$. Because, by [MV20, Lemma 3], $\mathbb{P}(\cup_{k\geq 0}\{\sigma_{k}=\infty\})=1$, this is indeed enough to imply the conclusion of [MV20, Lemma 6].
  • Assumption (A1) is also used in the proof of [MV20, Lemma 7]: the authors use it to prove that, for any $\alpha\in\mathcal{P}_{C}(E)$, $\mathbb{P}_{\alpha}(X(t)\in E)\geq\mathrm{e}^{-(1-c_{0})t}$ for some $c_{0}>\theta$ (in [MV20, Lemma 7], $c_{0}=c$ but this is not necessary). This is now given by Assumption (A’3) and thus the proof of [MV20, Lemma 7] goes through under our modified set of assumptions: one only needs to replace every occurrence of $c_{1}$ by $c_{0}$.
  • Finally, Assumption (A1) is used in the proof of Proposition 7 (Step 1 of the proof): the authors use the fact that $m_{n}P(E)\geq c_{1}n$ (=cn=cn in this paper) almost surely for all $n\geq 1$. Instead, we can use the fact that $m_{n}P(E)\geq cn/2$ almost surely for all $n$ large enough.
@@
Under Assumptions (T), (A’1), (A’2), (A’3), and (A4), if $m_{0}\cdot V<\infty$ and $m_{0}P\cdot V<\infty$, then the sequence of random measures $(m_{n}/n)_{n\geq 0}$ converges almost surely to $\nu R$ with respect to the topology of weak convergence. Furthermore, if $\nu R(E)>0$, then $(\tilde{m}_{n}:=m_{n}/m_{n}(E))_{n\in\mathbb{N}}$ converges almost surely to $\nu R/\nu R(E)$ with respect to the topology of weak convergence.
@@
We also let $\tilde{\eta}_{0}=0$ and, for all $n\geq 1$, $\tilde{\eta}_{n}=\frac{\eta_{n}}{n}$. The proof of [MV20] relies on the fact that $(\tilde{\eta}_{n})_{n\geq 0}$ is a stochastic approximation: indeed, for all $n\geq 0$, one can write
@@
where $F(\mu)=\mu Q-\mu Q(E)\mu$, $U_{n+1}=\tilde{\eta}_{n}Q(E)\delta_{\xi(n+1)}-\tilde{\eta}_{n}Q$, and $\gamma_{n}=\frac{1}{n\tilde{\eta}_{n-1}Q(E)}$.
We now introduce some continuous-time interpolation of $(\tilde{\eta}_{n})_{n\geq 0}$: For all $n\geq 0$, we set $\tau_{n}=\gamma_{1}+\cdots+\gamma_{n}$ and, for all $n\geq 0$, for all $t\in[\tau_{n},\tau_{n+1}]$,
$\tilde{\mu}_{t}=\tilde{\eta}_{n}+\frac{t-\tau_{n}}{\tau_{n+1}-\tau_{n}}(\tilde{\eta}_{n+1}-\tilde{\eta}_{n}).$
Remark 12 in [MV20] states that, without Assumption (A3), one can still conclude that $(\tilde{\mu}_{t})_{t\geq 0}$ is an asymptotic pseudo-trajectory in $\mathcal{P}_{C}(E)$ for the semi-flow induced in $\mathcal{P}_{C}(E)$ by the well-posed dynamical system
@@
(for all bounded functions $f=E\mapsto\mathbb{R}$). Now, by Theorem 3.7 in [Ben06], the limit set of a pre-compact asymptotic pseudo-trajectory is internally chain transitive and thus (by [Ben06, Proposition 5.3]) it does not contain any proper attractor. Now let $(X_{t})_{t\geq 0}$ be the continuous-time pure jump Markov process with sub-Markovian jump kernel QIQ-I; if, for all $t\geq 0$, we let $\mu_{t}$ be the distribution of $X_{t}$ conditioned on having survived until time $t$, then $(\mu_{t})_{t\geq 0}$ is a solution of (5.3). By Assumption (A’3), we get that $\nu$ is an attractor for the flow defined by (5.3) on $\mathcal{P}_{C}(E)$. Also, all trajectories started in $\mathcal{P}_{C}(E)$ converge to $\nu$. Thus, the limit set of $(\tilde{\eta}_{t})_{t\geq 0}$ is $\{\nu\}$. Therefore, $\tilde{\eta}_{n}\to\nu$ almost surely as $n\uparrow\infty$, for the topology of weak convergence. One can then follow the rest of the arguments in [MV20] to conclude the proof. ∎
5.2 Checking Assumption (A’1)